OpenStudy (anonymous):

Use the substitution method to find the solution(s) to the system of equations. Show all work for full credit. y = –x2 y = x

OpenStudy (anonymous):

\[y=x\] \[y=-x^2\]?

OpenStudy (anonymous):

that's what was on the question and it is one of the reasons i dont understand it

OpenStudy (anonymous):

if so set them equal to get \[-x^2=x\] \[x^2+x=0\] \[x(x+1)=0\] \[x=0\] or \[x=-1\]

OpenStudy (anonymous):

those are the two solutions and the check is easy

OpenStudy (anonymous):

thank you so much!!! :)

OpenStudy (anonymous):

yw